Stock Investment Transactions On September 12, 2,500 shares of Aspen Company are acquired at a price of $39.00 per share plus a $125 brokerage commission. On October 15, a $0.80-per-share dividend was received on the Aspen Company stock. On November 10, 1,000.00 shares of the Aspen Company stock were sold for $34 per share less a $50 brokerage commission. When required, round final answers to the nearest dollar. For a compound transaction, if an amount box does not require an entry, leave it blank. Prepare the journal entries for the original purchase, the dividend, and the sale under the cost method.

8/12

10/15

11/10
Answer #1
Sept 12 Investments-Aspen Company Stock 97625 =(2500*39)+125
         Cash 97625
Oct 15 Cash 2000 =2500*0.80
      Dividend Revenue 2000
Nov 10 Cash 33950 =(1000*34)-50
Loss on sale of Investments 5100
      Investments-Aspen Company Stock 39050 =97625*1000/2500